OpenStudy (anonymous):

ok, here is a good one for you... Given F(x)=sqrt(x+2) and g(x)=2x+1 determine (F-G)(x)

OpenStudy (anonymous):

When I did F(g(x)) I got sqrt(2x+3)

OpenStudy (anonymous):

(F-G)(x)=F(x)-G(x)=sqrt(x+2)-2x-1

OpenStudy (anonymous):

the other way around I am sitting at \[2(\sqrt{x+2})+1\]

OpenStudy (anonymous):

so am I doing this right then?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Why did you do F(g(x))?

OpenStudy (anonymous):

because I thought you had to substitute the values of G into the F function and then reverse it. I know I did that during my homework (this question is on my mid-term).

OpenStudy (anonymous):

If you have (FoG)(x) but if you have (F-G)(x) its different. Why reverse it? o.o

OpenStudy (anonymous):

yeah yeah.... the first one, I just wrote it wrong, sorry!

OpenStudy (anonymous):

so am I doing this properly?

OpenStudy (anonymous):

For FoG yes :)

OpenStudy (anonymous):

The first one sqrt(2x+3)

OpenStudy (anonymous):

sweet, thanks!

OpenStudy (anonymous):

can that second one be taken down further?

OpenStudy (anonymous):

You could make it sqrt(4x+8)+1. (just putting the 2 back in the sqrt) but other than that no. I would leave the 2 out of it.

OpenStudy (anonymous):

ok, cool... you rock, thanks!

OpenStudy (anonymous):

No problem :P